40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 3 18:49:44 2013, Gautam, Update, General, ITMx Oplev Fixed  
    Reply  Wed Jun 5 20:45:19 2013, gautam, Update, General, ITMx Oplev Fixed  ITMx_overview.JPGITMx_clipping_.JPG
       Reply  Wed Jun 26 19:02:38 2013, gautam, Update, General, ITMx Oplev (not quite?) Fixed oplev_chamber.pdfphoto_1.JPGphoto_2.JPG
          Reply  Thu Jun 27 18:11:53 2013, gautam, Update, General, ITMx Oplev-POX looks beam okay 
             Reply  Thu Jun 27 23:01:39 2013, gautam, Update, General, ITMx Oplev-servo gains adjusted pitch-plot.pdfyaw-plot.pdf
                Reply  Fri Jun 28 13:10:28 2013, gautam, Update, General, ITMx Oplev-servo gains adjusted pitch-plot_copy.pdfyaw-plot_copy.pdf
Message ID: 8758     Entry time: Wed Jun 26 19:02:38 2013     In reply to: 8684     Reply to this: 8770
Author: gautam 
Type: Update 
Category: General 
Subject: ITMx Oplev (not quite?) Fixed 

 

Summary:

Steve and I tried to fix the Oplev situation detailed in elog 8684, today afternoon. We have come up with a fix which needs to be adjusted, possibly completely overhauled depending on whether the mirror steering the return beam to the QPD is blocking the POX beam coming out. 

Details:

  • ITMx Oplev servo was first turned off.
  • It turns out we were hitting the wrong pair of Oplev steering mirrors inside the chamber. The incident beam was hitting a mirror meant for IR light (see sketch below) and not the intended first steering mirror. We pretty much redid the entire alignment (we used a pair of irises initially to set up a reference path) so as to hit the right pair of mirrors inside the chamber. At the end of today's efforts, the beam is reasonably well centered on both the intended mirrors, and there is not as much scattered light. 

Situation in the chamber: the black line is meant to indicate what was happening, the red is indicative of the present path.

 

oplev_chamber.pdf

  • The Oplev laser was installed in 2011 (October 13,2011 to be precise), and the quality of the beam coming out of the laser was pretty bad (the cross section was badly distorted even before it hit any optics). Steve thought this laser had reached the end of its lifetime, so we replaced the laser with a new one. Output power of this new laser has yet to be measured. The power-supply for the laser was also dodgy so we switched that out as well, and installed a new power supply. New laser and power supply are working satisfactorily. The old power supply will be checked with another laser tomorrow to gauge its status.
  • The cross-section of the beam from the new Oplev laser was deemed satisfactorily circular and we centred it on the first of the two lenses in the beam-path. We are getting 2.81 mW of power from the new laser.
  • In order to hit the right pair of Oplev steering mirrors while avoiding the clipping the beam on the tip-tilt/PR2 suspension, we had to sort of widen the angle of the beam going in, and so both steering mirrors, as well as the second lens in the beampath were shifted around. I have attached before and after pictures of the layout on the table. We adjusted things till the beam is reasonably well centred on both steering mirrors inside the chamber.
  • Because of the changed beam-path, the return Oplev beampath also changed, and so we had to move the steering mirror directing the return beam to the QPD as well. In its new position, it may be clipping the POX beam.
  • The beam is not getting clipped anywhere on the table now. It is also well centered on the two lenses in its path.
  • We placed two irises marking the new path so that we have a reference if the alignment needs to be changed again.
  • Turned ITMx Oplev servo back on.

Other stuff:

  1. The higher--power new laser means that the QPD sum is now ~6000counts (up from ~3500). The power in the return beam is 143.5 uW, which is ~5% of the power of the input beam.
  2. We centred the spot on the QPD, though when we excited ITM in yaw, the spot doesn't quite move horizontally, rather, it moves somewhat diagonally
  3. I turned the lights off, blocked the beam and measured the zero-current counts for the various QPD quadrants. These were all less than 1.5, and I reset these values on the ITMx Oplev screen according to my measurements.
  4. While viewing the spot of the Oplev beam on the ITMx face, we noticed that there were two spots, one less bright than the other. Steve suspects that this is because of multiple reflections from the chamber window.
  5. The status of the POX beam has to be verified (i.e is it getting clipped by the steering mirror for the return beam?)
  6. There was a Thorlabs PD on the table which had a green power LED. Jenne had asked me to cover this LED up, which I did with bits of tape.

Plan of action:

  • The spot size at the QPD is right now about 3mm. We may want to improve this by moving the first of the two lenses in the beam-path (there is not a whole lot of room to maneuvering the second lens.
  • Jenne just aligned and locked the X-arm, and doesn't think that the POX beam is being blocked, but I will verify this sometime tomorrow.

 

BEFORE

photo_1.JPG

 

AFTER

photo_2.JPG 

 

ELOG V3.1.3-